OpenStudy (anonymous):

intergration of x^3 e^x^2?

OpenStudy (anonymous):

come on sa-yui~you can do it~<support>! LOL

OpenStudy (anonymous):

lol., put x^2 = u then it will reduce to easy integration by parts :)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

so do you just use u=x^2 for integration and then substitute again?

OpenStudy (anonymous):

first tell me what u get after substituting u=x^2 ?

OpenStudy (anonymous):

x^3 e^u?

OpenStudy (anonymous):

when you substitute, you substitute entirely! there should be no traces of 'x' left so, you forst write x^3 = x* x^2 also you will HAVE TO also find, dx=.... ? x^2 =u so, du=... ?

OpenStudy (anonymous):

I got 1/2, integral of ue^u du?

OpenStudy (anonymous):

i get the same :) now isn't this the simplest integration by parts problem ? :)

OpenStudy (anonymous):

:D do you substitute u=x^2 first? or integrate u first?><

OpenStudy (anonymous):

we have already substituted u=x^2 to get 0.5* integral u e^u du now its time to integrate.

OpenStudy (anonymous):

isn't it 1/2 (ue^u)? :S

OpenStudy (anonymous):

answer after integration ? no0... thats what you are integrating... do you know what is integration by parts ?

OpenStudy (anonymous):

wait I'll try....><

OpenStudy (anonymous):

ok, take your time :)

OpenStudy (anonymous):

:O it doesn't work;( how can you integrate ue^u? :/ sorry:/

OpenStudy (anonymous):

what formula of integration by parts do you have ? (actually there are 2,i want to know which one is taught to u)

OpenStudy (anonymous):

oh! I got it now!:) thank you so much;)) I use the uv-integral v*dv/dx :)

OpenStudy (anonymous):

v*du/dx I mean:)

OpenStudy (anonymous):

oh, great! :) if you want me to verify, you can tell me what you got... oh, and since you are new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) if you have any doubts browsing this site, you can ask me, or to an ambassador (with an encircled A before their name). Have fun learning with Open Study! :)

OpenStudy (anonymous):

1/2 e^x^2 (x^2 -1):) thank you;)

OpenStudy (anonymous):

correct! good ;) welcome ^_^